OpenStudy (anonymous):

Pies are ordered from the Pizza Palace according to their diameter. Large:sixteen inches Medium:fourteen inches Small:twelve inches How many more square inches are in the large pizza than in the small pizza?

OpenStudy (anonymous):

\[\frac{\pi 8^2}{\pi 6^2}=\frac{16}{9}=1.77778 \]or\[\pi 8^2-\pi 6^2=28 \pi =88 \text{ sq inches} \]